# Combining Multi-Objective Bayesian Optimization with Reinforcement Learning for TinyML

### v2

@@ Abstract.
Deploying Deep Neural Networks (DNNs) on microcontrollers (TinyML) is a common trend to process the increasing amount of sensor data generated at the edge, but in practice, resource and latency constraints make it difficult to find optimal DNN candidates. Neural Architecture Search (NAS) is an excellent approach to automate this search and can easily be combined with DNN compression techniques commonly used in TinyML. However, many NAS techniques are not only computationally expensive, especially hyperparameter optimization (HPO), but also often focus on optimizing only a single objective, e.g., maximizing accuracy, without considering additional objectives such as memory consumption or computational complexity of a DNN, which are key to making deployment at the edge feasible. In this paper, we propose a novel NAS strategy for TinyML based on Multi-Objective Bayesian optimization (MOBOpt) and an ensemble of competing parametric policies trained using Augmented Random Search (ARS) Reinforcement Learning (RL) agents. Our methodology aims at efficiently finding tradeoffs between a DNN’s predictive accuracy, memory consumption on a given target system, and computational complexity. Our experiments show that we outperform existing MOBOpt approaches consistently on different data sets and architectures such as ResNet-18 and MobileNetV3.
Neural Architecture Search, TinyML, Reinforcement Learning, Multi-Objective Optimization
copyright: acmlicensedjournalyear: 2023doi: XXXXXXX.XXXXXXXbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning algorithmsccs: Computing methodologies Search methodologiesccs: Computing methodologies Reinforcement learning
@@ 1. Introduction
The deployment of DNNs on embedded devices is restricted to the constraints imposed by the target microcontroller platform. Constraints such as $<$1MB Flash, $<$512Kb RAM, and clock speeds in the low MHz range make the design of DNN models for such platforms challenging. In general, several conflicting objectives and constraints such as memory availability, inference time, and power consumption of the deployed DNN model must be considered.
@@
Despite extensive research on Neural Architecture Search (NAS), there is still no definitive method that is both fast and robust. Many black-box optimization approaches (Zoph and Le, 2017; Real et al., 2019; White et al., 2021) work reliably, but are inefficient. Differentiable NAS strategies such as DARTS (Liu et al., 2019) have robustness problems, as it can be observed that the architectures found often do not generalize well because the search overfits the validation data set (Zela et al., 2020). Finally, zero-cost NAS strategies that have been proposed to allow rapid network specialization for different target platforms and resource constraints, such as Once-for-all (Cai et al., 2020) or PreNAS Wang et al., 2023, while fast, do not provide accurate information about an architecture’s performance, but only statistics, such as the number of FLOPs (White et al., 2022).
@@
In this work, we investigate how to improve the efficiency of NAS for DNN deployment on microcontrollers using a combination of Multi-objective Bayesian optimization (MOBOpt), deep compression, and Augmented Random Search (ARS) (Mania et al., 2018). To this end, we formulate the architecture search as a hyperparameter optimization (HPO) problem (Bergstra et al., 2013), where the parameters influence the complexity of a given architecture using filter pruning (Li et al., 2017) and weight quantization (Gholami et al., 2021). This allows us to optimize DNNs, considering the objectives of top-1 accuracy, memory consumption, and FLOPS.
@@
As our main contribution, we propose a novel Bayesian-optimization based solver using an ensemble of competing locally parameterizable policies that are iteratively trained on the underlying surrogate model using an ARS RL agent. Our experiments show that this RL-based strategy helps to improve the effectiveness of our NAS approach in finding optimal DNNs for microcontroller deployment compared to other state-of-the-art multi-objective optimization strategies for multiple use cases while introducing negligible additional overhead for training the RL agents. The architectures proposed by our approach can be deployed directly on common microcontrollers without further (re-)training.
@@
The rest of the paper is structured as follows. Sec. 2 discusses related work. Sec. 3 provides background on multi-objective Bayesian optimization. Sec. 4 describes our RL agent and how we integrate it into our MOBOpt DNN hyperparameter exploration process. We evaluate our approach in Sec. 5 on several public datasets with different types of neural networks (Sec. 5.2) and on a synthetic optimization problem (Sec. 5.4). Sec. 6 concludes.
@@ 2. Related Work
Different approaches for efficient design and deployment of DNNs on embedded platforms have been prominently discussed in scientific research: Dedicated architectures such as the MobileNet class (Howard et al., 2017; Sandler et al., 2018; Howard et al., 2019) have been proposed, introducing scaling parameters and specialized layers, i.e., depthwise-separable convolutions, to control size and inference time. Deep Compression features a number of techniques, including pruning (Li et al., 2017), which dynamically removes weights during training, and weight quantization (Jacob et al., 2018), which reduces the resolution at which weights are represented. Knowledge distillation (KD) (Hinton et al., 2015) has also been used for DNN compression (Ashok et al., 2018; Cao et al., 2018). However, DNN optimization using KD remains challenging as the student often fails to match the predictions of the teacher (Stanton et al., 2021; Cho and Hariharan, 2019). Based on DNN compression deployment pipelines for embedded devices have been proposed: (Han et al., 2016) combine network pruning, weight quantization and Huffman coding and follow-up work compares different pruning and quantization methods (Deutel et al., 2023). Other well known DNN deployment pipelines are MCUNet (Lin et al., 2020, 2021) and Tensorflow Lite Micro (David et al., 2021).
@@
Although NAS is an intensively researched topic, harnessing its power to find DNN deployment candidates for microcontroller platforms is still a largely unsolved problem. Previous approaches (Wang et al., 2023; Cai et al., 2020, 2018) mainly focus on zero-cost NAS that derive architecture variants for a given target platform and a given set of resource constraints from a well-trained ”supernet” structure or focus on hardware aware search for larger mobile platforms (Tan et al., 2019; Dong et al., 2018). In contrast, although largely unexplored in the context of NAS for microcontroller targets, Multi-Objective Bayesian Optimization (MOBOpt) is one of the most widely used black-box optimization methods in NAS for solving HPO problems. (Elsken et al., 2018; White et al., 2021; Chitty-Venkata and Somani, 2022; White et al., 2023). As training DNNs is time- and resource-intense the reduced sample complexity of Bayesian optimization is beneficial compared to otherwise popular genetic optimization (Lu et al., 2019; Deb et al., 2002). When optimizing continuous hyperparameters, algorithms using Gaussian processes (GP) are most prominent (Knowles, 2006). An extension of this optimization approach that specifically aims to tackle problems with high-dimensional search spaces suggests the use of locally modeled confidence regions (Eriksson et al., 2019; Daulton et al., 2022).
The use of RL for MOOpt has been extensively studied: The general approach is to consider the search space as the action space of the RL agent, whose reward is the performance of the trained architecture on unseen data. Approaches to optimization using RL differ in how they represent the policies they use, e.g., (He et al., 2018) use Deep Deterministic Policy Gradient (DDPG) agents, (Moffaert and Nowé, 2014) use tabular Q-learning, (Kim et al., 2022) use single-step RL approaches, and (Li et al., 2020) utilize actor-critic algorithms.
@@
In this work, we combine the sampling efficiency of GPs with the expressive power of RL agents. Compared to previous work on RL for MOOpt, our approach does not directly apply RL to the search space of the MOOpt problem, but instead uses RL to improve the optimization performed on the GPs. To address the problem of expending large amounts of resources to train RL agents, we instead propose to train a set of competing policies within MOBOpt that implement candidate sampling using very simple Augmented Random Search (ARS) (Mania et al., 2018) RL agents. Hence, our approach allows for efficient exploration of DNN architectures directly deployable on microcontroller.
@@ 3. Background
Hence, the goal of MOOpt is to identify an approximate PF $\mathcal{P}(X_{n})$ of the true PF $\mathcal{P}(X)$ within a search budget of $n$ evaluations. In cases where the true PF is not known, the quality of $\mathcal{P}(X_{n})$ is usually assessed using the Hypervolume (HV) indicator, also called the S-metric.
@@
A feasible reference point is commonly derived from domain knowledge of the problem. The HV also allows to compare multiple PFs as long as they were calculated with the same reference point and share the same objective values.
A way of improving the sample efficiency of multi-objective optimization is by using Bayesian Optimization (BO) aiming at minimizing the number of required evaluations of the given parameter space. BO treats the objective functions as black-boxes and places a prior over them, thereby capturing beliefs about their behaviour. Over time, as new samples are collected, the prior is updated iteratively to form the posterior distribution over the objective functions, in our case by using GPs. The posterior, in turn, is used in each iteration to determine the next sample-point to evaluate by employing an acquisition function as a heuristic to quantize the ”usefulness” of the sample. Traditionally, the acquisition function is also responsible for balancing the exploration and exploitation tradeoff. Furthermore, since we are considering multiple objectives, we scale and accumulate the objective values to form a single-objective problem (Knowles, 2006; Paria et al., 2020).
A commonly used acquisition function is the Expected Improvement (EI) (Močkus, 1975)
@@
where $f^{*}$ is the objective value of the best observed sample so far. For a given Bayesian model the EI can be evaluated with an integral over the posterior distribution either analytically or by Monte-Carlo sampling (Wilson et al., 2017), which allows to approximate the EI as
@@
  Parameters: search budged $J$, learning rate $\alpha$, directions sampled per iteration $N$, rollout horizon $H$, constant exploration noise $v$, number of top-performing directions to use $k$, number of objectives $n$, number of parameters $p$, number of ARS agents $L$, objective function $f(x)=[f_{1}(x),\dots,f_{n}(x)]$
@@
     Select $\{x_{1},x_{2},\dots,x_{L}\}$ initial states from $\mathcal{P}_{feas}$ of previous evaluations using k-means clustering.
@@
           Update the parameters $\theta_{l}$ of policy $\pi_{l}$ using the top-$k$ performing rollouts (with $\theta_{l,0}=0$)
@@ 4. Method
In this work we solve a multi-objective HPO problem using Bayesian optimization and RL. Different from other Bayesian optimization algorithms, our methodology iteratively finds new sets of hyperparameters using an ensemble of competing multi-layer perceptron (MLP)-based RL agents, in particular ARS (Mania et al., 2018), to optimize the acquisition function. Our policies only use two linear layers with a hidden layer size of 64 resulting in a couple of thousand trainable parameters each. ARS is an RL agent which learns a set of linear (or MLP) policies $\pi_{\theta}$, parameterized by a set of vectors $\theta$, one for each layer, for controlling a dynamic environment $\mathbb{E}_{\xi}$, with $\xi$ encoding the randomness of the environment, maximizing/minimizing an average reward/loss $r(\pi_{\theta},\xi)$. ARS achieves this by optimizing over the set of parameters $\theta$ by utilizing derivative-free optimization with noisy function evaluations and for each layer iteratively performing updates using directions of the form $\frac{r(\pi_{\theta+\nu\varphi},\xi_{1})-r(\pi_{\theta-\nu\varphi},\xi_{2})}{\nu}$ for two i.i.d. random variables $\xi_{1}$ and $\xi_{2}$, $\nu$ a positive real number ($0.008$ in our case), and $\varphi$ a zero mean Gaussian vector with i.i.d. standard normal entries. The two rewards $r(\pi_{\theta+\nu\varphi},\xi_{1})$ and $r(\pi_{\theta-\nu\varphi},\xi_{2})$ are obtained by collecting two sets of trajectory rollouts from the system of interest.
@@
We initialize the training of our competing ARS agents with different samples from the current feasible Pareto front $\mathcal{P}_{\text{{feas}}}(X)$. To find the most diverse set of samples, we perform k-means clustering on the Pareto front if it consists of more elements than we can train policies. To query the objective values, i.e., accuracy, memory consumption, and FLOPS, for a proposed set of hyperparameters of a given DNN architecture, our objective function evaluator first trains, prunes, and quantizes the DNN. Afterwards, the DNN is converted into source code using an automatic code generator to accurately assess memory usage of weights and activation tensors. Our implementation performs this procedure using the compression and deployment pipeline proposed by (Deutel et al., 2023), who describe a configurable pipeline that supports both DNN pruning and full weight quantization. Following their results, we focus our experiments on iterative filter pruning and post-training static quantization.
@@
We integrated ARS into MOBOpt as a parameterized, trainable solver, which can thus be seen as an alternative to traditional Bayesian solvers, see Algorithm 1. The derivative-free random sampling approach implemented by ARS is computationally inexpensive and highly parallelizable, allowing for efficient GPU-accelerated execution.
@@
We define the states $x_{0},x_{1},\dots,x_{L}$ of our $L$ competing agents as vectors of the parameters of the search space of our optimization problem. Each agent performs actions based on its trainable policy $\pi_{\theta}$ by altering the parameter values of its state vector within given bounds. We train the competing policies of our ARS agents for each sample after evaluating the objective function by performing rollouts on posteriors sampled from the GP. We fit them by maximizing the marginal likelihood of the previously queried evaluation of the objective function (see also (Williams and Rasmussen, 1995)). The policy training performed for each sample ends after a maximum number of training steps has been conducted. To obtain the initial prior for the GP, we propose candidates during the first couple of iterations of the optimization using Latin Hypercube sampling (McKay et al., 2000). We chose this approach over random sampling because we found that by using Latin Hypercube sampling, we could significantly improve the expressiveness of our priors and therefore enhance the performance of all evaluated GP-based sampling strategies.
@@
As a reward $r(x)$, we first scale our $n=4$ objectives (accuracy, ROM/RAM memory consumption, FLOPS) to a single objective value using augmented Chebyshev scalarization (Knowles, 2006), see Eq. (7), where $\rho=0.005$ is a small positive value and $\lambda$ is a weight vector that we draw uniformly at random at the beginning of each sample. We then calculate the difference between the scaled objective and the best previously visited (scaled) objective $f^{*}$ as the reward, see Eq. (8).
@@
As a result, the reward $r(x)$ used by our ARS agents is comparable to acquisition functions used in traditional Bayesian optimization. In particular, our approach is similar to the expected improvement (EI) used by ParEgo (Knowles, 2006), with the notable difference that we consider penalized ”negative improvements” instead of clamping them to zero. The main reason for clamping ”negative improvement” in combination with a Monte Carlo-based evaluation of the GPs is to increase the attractiveness of regions with high uncertainty, and is therefore the mechanism of EI to control the tradeoff between exploration and exploitation. However, we argue that we can still achieve the desired tradeoff if we instead penalize ”negative improvements” by multiplication with a small constant factor, while allowing our ARS policies to learn from situations where the current objective function evaluation is worse than the best previously visited one.
After training the competing ARS agents, we perform several rollouts for each of them starting from their respective starting points selected from $\mathcal{P}_{\text{{feas}}}(X)$. We then select the policy that yielded the overall best reward to propose the next set of hyperparameters to be evaluated on the objective function in the next sample. A variant of this approach would be to consider a set of best-performing parametrizations as candidates for a batched optimization in different segments of the Pareto front.

### v3

@@ Abstract.
Deploying deep neural networks (DNNs) on microcontrollers (TinyML) is a common trend to process the increasing amount of sensor data generated at the edge, but in practice, resource and latency constraints make it difficult to find optimal DNN candidates. Neural architecture search (NAS) is an excellent approach to automate this search and can easily be combined with DNN compression techniques commonly used in TinyML. However, many NAS techniques are not only computationally expensive, especially hyperparameter optimization (HPO), but also often focus on optimizing only a single objective, e.g., maximizing accuracy, without considering additional objectives such as memory requirements or computational complexity of a DNN, which are key to making deployment at the edge feasible. In this paper, we propose a novel NAS strategy for TinyML based on multi-objective Bayesian optimization (MOBOpt) and an ensemble of competing parametric policies trained using Augmented Random Search (ARS) reinforcement learning (RL) agents. Our methodology aims at efficiently finding tradeoffs between a DNN’s predictive accuracy, memory requirements on a given target system, and computational complexity. Our experiments show that we consistently outperform existing MOBOpt approaches on different datasets and architectures such as ResNet-18 and MobileNetV3.
Neural Architecture Search, Embedded Systems, Reinforcement Learning, Multi-Objective Bayesian Optimization
copyright: acmlicensedjournal: TELOjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3715012booktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Reinforcement learningccs: Computer systems organization Embedded softwareccs: Computing methodologies Continuous space search
@@ 1. Introduction
The deployment of DNNs on embedded devices is restricted to the constraints imposed by the target microcontroller platform. Constraints such as $<$1MB Flash, $<$512Kb RAM, and clock speeds in the low-MHz range make the design of DNN models for such platforms challenging. In general, several conflicting objectives and constraints such as memory availability, inference time, and power of the deployed DNN model must be considered.
@@
Despite extensive research on neural architecture search (NAS), there is still no definitive method that is both fast and robust. Many black-box optimization approaches (Zoph and Le, 2017; Real et al., 2019; White et al., 2021) work reliably, but are inefficient. Differentiable NAS strategies such as DARTS (Liu et al., 2019) have robustness problems, as it can be seen that the architectures found often do not generalize well because the search overfits the validation dataset (Zela et al., 2020). Finally, zero-cost NAS strategies that have been proposed to allow rapid network specialization for different target platforms and resource constraints, such as Once-for-all (Cai et al., 2020) or PreNAS Wang et al., 2023, while fast, do not provide accurate information about an architecture’s performance, but only statistics, such as the number of FLOPs (White et al., 2022).
@@
In this work, we investigate how to improve the efficiency of NAS for DNN deployment on microcontrollers using a combination of multi-objective Bayesian optimization (MOBOpt), deep compression, and Augmented Random Search (ARS) (Mania et al., 2018). To this end, we formulate the architecture search as a hyperparameter optimization (HPO) problem (Bergstra et al., 2013), with the hyperparameters controlling the use of filter pruning (Li et al., 2017) and weight quantization (Gholami et al., 2022). As optimization objectives, we consider top-1 accuracy, memory (ROM and RAM), and FLOPS.
Our proposed HPO method consists of two building blocks: (1) the multi-objective optimization loop and (2) the objective functions evaluator, see Fig. 1. Our proposed search strategy iteratively determines new sets of training, pruning, and quantization hyperparameters until a termination condition is met. The trained DNN resulting from the use of each of the proposed sets of hyperparameters is then evaluated, resulting in a vector of scalar objective values.
Figure 1. Overview of our proposed RL-based multi-objective Bayesian DNN hyperparameter optimization approach for microcontroller targets.
[…]
Our multi-objective DNN hyperparameter optimization approach for microcontrollers uses Augmented Random Search RL agents to iteratively propose new DNN configurations, which in turn are evaluated using objective function evaluators for the four considered objectives: FLOPs, accuracy, ROM, and RAM.
@@
Our main contribution, a novel Bayesian-optimization based solver, uses an ensemble of competing locally parameterizable policies that are iteratively trained on the underlying surrogate model using ARS RL agents. Our experiments show that this RL-based strategy helps to improve the effectiveness of our NAS approach in finding optimal DNNs for microcontroller deployment compared to other state-of-the-art multi-objective optimization strategies for multiple use cases while introducing negligible additional overhead for training the RL agents. The architectures proposed by our approach can be deployed directly on common microcontrollers without further (re-)training. To make the results presented in this work accessible, we have open-sourced the implementation of our method [^1] A Python implementation of the Bayesian optimization-based solver discussed in this work, including the implementation of both the ARS and the PPO RL agents, is available at https://doi.org/10.5281/zenodo.13918411. The repository includes documentation on how to install and execute the code..
@@
The remainder of the paper is structured as follows. Section 2 discusses related work. Section 3 provides background on multi-objective Bayesian optimization. Section 4 describes our RL agents and how we integrate them into our MOBOpt DNN hyperparameter exploration process. We evaluate our approach in Section 5. First, we compare our MOBOpt approach with other optimizers (Bayesian and non-Bayesian) on several public datasets and with two types of DNN, MobileNetV3 and ResNet18, see Section 5.1. Second, in Section 5.2 and 5.3, we evaluate the robustness and performance of the ARS agents by discussing the hyperparameter selection and results on a small synthetic optimization problem. Third, we compare ARS with PPO, an alternative deep-RL algorithm, and compare our approach to other optimization strategies such as Automated Pruning and Zero-Shot NAS in Section 5.4.
@@ 2. Related Work
Different approaches for efficient design and deployment of DNNs on embedded platforms have been prominently discussed in scientific research: Dedicated architectures such as the MobileNet class (Howard et al., 2017; Sandler et al., 2018; Howard et al., 2019) have been proposed, introducing scaling parameters and specialized layers, i.e., depthwise-separable convolutions, to control size and inference time. Deep compression features a number of techniques, including pruning (Li et al., 2017), which dynamically removes weights during training, and weight quantization (Jacob et al., 2018), which reduces the resolution at which weights are represented. Knowledge distillation (KD) (Hinton et al., 2015) has also been used for DNN compression (Ashok et al., 2018; Cao et al., 2018). However, DNN optimization using KD remains challenging, as the student often fails to match the predictions of the teacher (Stanton et al., 2021; Cho and Hariharan, 2019). Based on DNN compression deployment pipelines for embedded devices have been proposed: Han et al. (2016) combine network pruning, weight quantization, and Huffman coding, and follow-up work compares different pruning and quantization methods (Deutel et al., 2023). Other well-known DNN deployment pipelines are MCUNet (Lin et al., 2020, 2021) and Tensorflow Lite Micro (David et al., 2021).
@@
Although NAS is an intensively researched topic, harnessing its power to find DNN deployment candidates for microcontroller platforms is still a largely unsolved problem. Previous approaches (Wang et al., 2023; Cai et al., 2020, 2018) focus mainly on zero-cost NAS that derive architecture variants for a given target platform and a given set of resource constraints from a well-trained ”supernet” structure or focus on hardware-aware search for larger mobile platforms (Tan et al., 2019; Dong et al., 2018). In contrast, although largely unexplored in the context of NAS for microcontroller targets, multi-objective Bayesian Optimization (MOBOpt) is one of the most widely used black-box optimization methods in NAS to solve HPO problems. (Elsken et al., 2018; White et al., 2021; Chitty-Venkata and Somani, 2022; White et al., 2023). As DNN training is time- and resource-intensive the reduced sample complexity of Bayesian optimization is beneficial compared to otherwise popular genetic optimization (Lu et al., 2019; Deb et al., 2002). When optimizing continuous hyperparameters, algorithms using Gaussian processes (GP) are most prominent (Knowles, 2006). An extension of this optimization approach that specifically aims to tackle problems with high-dimensional search spaces suggests the use of locally modeled confidence regions (Eriksson et al., 2019; Daulton et al., 2022).
@@
The use of RL for MOOpt has been extensively studied: The general approach is to consider the search space as the action space of the RL agent, whose reward is the performance of the trained architecture on unseen data. Approaches to optimization that use RL differ in how they represent the policies they use, e.g., He et al. (2018) use Deep Deterministic Policy Gradient (DDPG) agents, Moffaert and Nowé (2014) use tabular Q-learning, Kim et al. (2022) use single-step RL approaches, and Li et al. (2020) utilize actor-critic algorithms. In this work, we combine the sampling efficiency of GPs with the expressive power of RL agents. Compared to previous work on RL for MOOpt, our approach does not directly apply RL to the search space of the MOOpt problem, but instead uses RL to improve the optimization performed on the GPs. To address the problem of expending large amounts of resources to train RL agents, instead we propose to train a set of competing policies within MOBOpt that implement candidate sampling using very simple Augmented Random Search (ARS) (Mania et al., 2018) RL agents. Hence, our approach allows for efficient exploration of DNN architectures that can be directly deployed on microcontrollers.
Besides MOOpt, RL has been used extensively for automatic DNN pruning as well. Pruning strategies such as AMC (He et al., 2018), DECORE (Alwani et al., 2022) or NEON (Hirsch and Katz, 2022) use RL to suggest sparsity ratios per layer for a base model to perform pruning. Either a single RL agent is applied iteratively layer by layer, suggesting sparsity ratios for each prunable layer individually (AMC and NEON), or a different RL agent is used for each layer, resulting in a multi-agent approach (DECORE). As a result, the RL agents replace the tuning of pruning hyperparameters usually performed by a human engineer. Such automatic pruning approaches differ significantly from our approach and from MOOpt in general in that they do not explore trade-offs between different pruning schedules and hyperparameter configurations while iteratively learning from successive separate DNN training and pruning processes, but instead automatically search for a sensible pruning configuration as part of a single run. As a result, automatic pruning does not produce a Pareto front like MOOpt would, nor does it consider other hyperparameters that influence training, such as training epochs, batch size, or learning rate, but is limited to exploring sparsity ratios alone.
As an alternative to offline pruning, which is performed as part of DNN training, dynamic pruning describes methods that select a subset of layer structures of a DNN at runtime based on the processed inputs (Wang et al., 2018; Gao et al., 2019; Liu and Deng, 2018; Bolukbasi et al., 2017). Compared to MOOpt as presented in this work, where a Pareto front of different and unique architectural tradeoffs between accuracy, memory, and computational complexity is searched offline, dynamic pruning never actually removes structures of an architecture completely, but instead dynamically selects different subsets based on the input it receives online. As a result, dynamic pruning only reduces the computational effort of DNN inference, but not the problem of limited memory (RAM, ROM) as considered here in the context of TinyML.
@@ 3. Background
Hence, the goal of MOOpt is to identify an approximate feasible PF $\mathcal{P}_{feas}(X_{n})$ of the true PF $\mathcal{P}(X)$ within a search budget of $n$ evaluations. In cases where the true PF is not known, the quality of $\mathcal{P}_{feas}(X_{n})$ is usually assessed using the Hypervolume (HV) indicator, also called the S-metric.
@@
A feasible reference point is commonly derived from the domain knowledge of the problem. The HV also allows one to compare multiple PFs as long as they were calculated with the same reference point and share the same objective values.
A way of improving the sample efficiency of multi-objective optimization is by using Bayesian Optimization (BO) aiming at minimizing the number of required evaluations of the given parameter space. BO treats the objective functions as black-boxes and places a prior over them, thereby capturing beliefs about their behavior. Over time, as new samples are collected, the prior is iteratively updated to form the posterior distribution over the objective functions, in our case using GPs. The posterior, in turn, is used in each iteration to determine the next sample-point to evaluate by employing an acquisition function as a heuristic to quantize the ”usefulness” of the sample. Traditionally, the acquisition function is also responsible for balancing the exploration and exploitation tradeoff. Furthermore, since we are considering multiple objectives, we scale and accumulate the objective values to form a single-objective problem (Knowles, 2006; Paria et al., 2020).
A common acquisition function is the Expected Improvement (EI) (Močkus, 1975), see Eq. (1),
@@
where $f^{*}$ is the objective value of the best observed sample so far. For a given Bayesian model the EI can be evaluated with an integral over the posterior distribution either analytically or by Monte-Carlo sampling (Wilson et al., 2017), which allows to approximate the EI as shown in Eq. (2),
@@
[…]
The Multi-Objective Bayesian Optimization with Augmented Random Search that we propose in this paper and that we describe in detail in the Methods section of this paper.
  Parameters: search budged $J$, learning rate $\alpha$, directions sampled per iteration $N$, rollout horizon $H$, constant exploration noise $v$, number of top-performing directions to use $b$, number of objectives $n$, number of parameters $p$, number of ARS agents $L$, objective function $f(x)=[f_{1}(x),\dots,f_{n}(x)]$
@@
     Select $\{x_{1},x_{2},\dots,x_{L}\}$ initial states from previous evaluations $\mathcal{P}_{feas}$ using k-means clustering.
@@
           Update the parameters $\theta_{l}$ of policy $\pi_{l}$ using the top-$b$ performing rollouts (with $\theta_{l,0}=0$)
@@ 4. Method
In this work, we solve a multi-objective HPO problem using a combination of Bayesian optimization and RL. Unlike other Bayesian optimization algorithms, our methodology iteratively finds new sets of hyperparameters using an ensemble of competing multi-layer perceptron (MLP)-based RL agents, in particular ARS (Mania et al., 2018), to optimize the acquisition function. The RL agents operate on the hyperparmaeters as their action space. Our policies only use two linear layers with a hidden layer size of 64 resulting in a couple of thousand trainable parameters each. ARS is an RL agent that learns a set of linear (or MLP) policies $\pi_{\theta}$, parameterized by a set of vectors $\theta$, one for each layer, to control a dynamic environment $\mathbb{E}_{\xi}$, with $\xi$ encoding the randomness of the environment, maximizing/minimizing an average reward/loss $r(\pi_{\theta},\xi)$. ARS achieves this by optimizing over the set of parameters $\theta$ by utilizing derivative-free optimization with noisy function evaluations and for each layer iteratively performing updates using directions of the form $\frac{r(\pi_{\theta+\nu\varphi},\xi_{1})-r(\pi_{\theta-\nu\varphi},\xi_{2})}{\nu}$ for two i.i.d. random variables $\xi_{1}$ and $\xi_{2}$, $\nu$ a positive real number ($0.008$ in our case), and $\varphi$ a zero mean Gaussian vector with i.i.d. standard normal entries. The two rewards $r(\pi_{\theta+\nu\varphi},\xi_{1})$ and $r(\pi_{\theta-\nu\varphi},\xi_{2})$ are obtained by collecting two sets of trajectory rollouts from the system of interest.
@@
We initialize the training of our competing ARS agents with different samples from the current feasible Pareto front $\mathcal{P}_{\text{{feas}}}(X)$. To find the most diverse set of samples, we perform k-means clustering on the Pareto front if it consists of more elements than we can train policies. To query the objective values, i.e., accuracy, memory, and FLOPS, for a proposed set of hyperparameters of a given DNN architecture, our objective function evaluator first trains, prunes, and quantizes the DNN. Afterwards, the DNN is converted into source code using an automatic code generator to accurately assess the memory usage of the weights and activation tensors. Our implementation performs this procedure using the compression and deployment pipeline proposed by Deutel et al. (2023), who describe a configurable pipeline that supports both DNN pruning and full weight quantization. Following their results, we focus our experiments on iterative filter pruning and post-training static quantization.
@@
We integrated ARS into MOBOpt as a parameterized trainable solver, which can thus be seen as an alternative to traditional Bayesian solvers, see Algorithm 1. The derivative-free random sampling approach implemented by ARS is computationally inexpensive and highly parallelizable, allowing efficient GPU-accelerated execution.
@@
We define the states $x_{0},x_{1},\dots,x_{L}$ of our $L$ competing agents as vectors of the parameters of the search space of our optimization problem. Each agent performs actions based on its trainable policy $\pi_{\theta}$ by altering the parameter values of its state vector within given bounds. We train the competing policies of our ARS agents for each sample after evaluating the objective function by performing rollouts on posteriors sampled from the GP. We fit them by maximizing the marginal likelihood of the previously queried evaluation of the objective function (see also Williams and Rasmussen (1995)). The policy training performed for each sample ends after a maximum number of training steps have been conducted. To obtain the initial prior for the GP, we propose candidates during the first ten iterations of optimization using Latin Hypercube sampling (McKay et al., 2000). We chose this approach over random sampling because we found that by using Latin Hypercube sampling, we could significantly improve the expressiveness of our priors and, therefore, enhance the performance of all evaluated GP-based sampling strategies.
@@
As a reward $r(x)$, we first scale our $n=4$ objectives (accuracy, ROM/RAM memory, FLOPS) to a single objective value using augmented Chebyshev scalarization (Knowles, 2006), see Eq. (7), where $\rho=0.005$ is a small positive value and $\lambda$ is a weight vector that we draw uniformly at random at the beginning of each sample. We then calculate the difference between the scaled objective and the best previously visited (scaled) objective $f^{*}$ as the reward, see Eq. (8).
@@
As a result, the reward $r(x)$ used by our ARS agents is comparable to the acquisition functions used in traditional Bayesian optimization. In particular, our approach is similar to the expected improvement (EI) used by ParEgo (Knowles, 2006), with the notable difference that we consider penalized ”negative improvements” instead of clamping them to zero. The main reason for clamping ”negative improvement” in combination with a Monte Carlo-based evaluation of the GPs is to increase the attractiveness of regions with high uncertainty, and is therefore the mechanism of EI to control the tradeoff between exploration and exploitation. However, we argue that we can still achieve the desired tradeoff if we instead penalize ”negative improvements” by multiplication with a small constant factor, while allowing our ARS policies to learn from situations where the current objective function evaluation is worse than the best previously visited one.
After training the competing ARS agents, we perform several rollouts for each of them starting from their respective starting points selected from $\mathcal{P}_{\text{{feas}}}(X)$. We then select the policy that yielded the overall best reward to propose the next set of hyperparameters to be evaluated on the objective function in the next sample. A variant of this approach would be to consider a set of best-performing parameterizations as candidates for a batched optimization in different segments of the Pareto front.